OpenStudy (anonymous):

Which of the following statements have the same result? Explain each step in solving each one. f(3) when f(x) = 2x + 2 f-1(4) when f(x) = 3x minus 4, all over 5 y + 10 = 2y + 1

OpenStudy (anonymous):

@SolomonZelman @SCARLET_SPIDER

OpenStudy (anonymous):

f(3) when f(x)=2x+2 f=2(3)+2 f=5+2 f=7

OpenStudy (anonymous):

f(4) when f(x)=(3x-5)/4 f=(3*4-5)/4 f=(12-5)/4 f=7/4 f=1.75

OpenStudy (anonymous):

Thanks :3

OpenStudy (anonymous):

y+10=2y+1 -1 -1 y+9=2y -y -y 9=y

OpenStudy (anonymous):

The other one, i'm not sure because you put f-1(4)

OpenStudy (anonymous):

No prob :)